# UniHR: Hierarchical Representation Learning for Unified Knowledge Graph Link Prediction

### v3

@@ 5.1 Experiment Settings
Evaluation Metric.
We conduct link prediction across multiple settings, evaluating performance based on the rank of predicted facts. We use the MR (Mean Rank), MRR (Mean Reciprocal Rank) and Hits@K (K=1,3,10) as our evaluation metrics. We abbreviate ‘Hits@K’ as ‘H@K’ and employ filtering settings [4] during the evaluation to eliminate existing facts in the dataset. It is worth noting that for the query $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ in HKGs, there are two evaluation filtering settings in existing models: one that filters out facts satisfying $((h,r,?),{(k_{i}:v_{i})}_{i=1}^{m})$ and another that filters out facts satisfying only $(h,r,?)$ in the training set. Similarly, the difference in filtering settings of TKG occurs in timestamp. In this paper, we adopt the strict filtering setting of the former. To ensure fair comparison, for HKG we utilize the results of HyNT [6] with the same settings as ours. For TKG, we thoroughly review the original code of our baselines and reproduce the results of some methods.
Baselines.
For link prediction on HKG, we compare our UniHR with NaLP [15], tNaLP [15], RAM [22], HINGE [27], NeuInfer [14], StarE [12], HyTransformer [43], GRAN [34], HyNT [6] and HAHE [23]. For link prediction on NKG, QuatE [45], BiQUE [16], Neural-LP [42], DRUM [28], AnyBURL [24], BiVE [7] and NestE [37] are chosen as baselines. Among them, BiVE and NestE are especially designed for NKG. We compare against following TKG link prediction methods: ComplEx-N3 [20], HyTE [8], TA-DistMult [13], ATiSE [38], TeRo [39], TGeomE+ [40], HGE [26] and ECEformer [11].
Implementation details.
All experiments are conducted on a single Nvidia 80G A800 GPU and implemented with PyTorch. For base link prediction on NKGs, we also use augmented triples from [7] for training to ensure fairness. For triple prediction, due to the small size of training set, we conduct training based on fixed embeddings of entities obtained from the base link prediction and set $\omega_{nested\;relation}\text{=}\ 0$ to prevent overfitting. We employ AdamW [19] optimizer. Hyperparameter details can be found in Appendix C.
5.2 Main Results
Table 2: Results of link prediction on HKG datasets. Results are reported by us, and the others are taken from [6]. Best results are bold, while the second are underlined.
WikiPeople WD50K
Model subject/object all entities subject/object all entities
MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10 MRR H@1 H@10
NaLP [15] 0.356 0.271 0.499 0.360 0.275 0.503 0.230 0.170 0.347 0.251 0.187 0.375
tNaLP [15] 0.358 0.288 0.486 0.361 0.290 0.490 0.221 0.163 0.331 0.243 0.182 0.360
RAM [22] 0.459 0.384 0.584 0.461 0.386 0.585 0.276 0.210 0.399 0.296 0.232 0.416
HINGE [27] 0.393 0.309 0.547 0.395 0.311 0.549 0.264 0.187 0.410 0.277 0.200 0.424
NeuInfer [14] 0.357 0.247 0.533 0.357 0.248 0.532 0.220 0.154 0.347 0.225 0.158 0.355
StarE [12] 0.458 0.364 0.611 - - - 0.309 0.234 0.452 - - -
HyTransformer [43] 0.465 0.395 0.598 - - - 0.308 0.239 0.452 - - -
GRAN [34] 0.477 0.408 0.596 0.480 0.411 0.599 0.329 0.259 0.465 0.363 0.294 0.493
HyNT [6] 0.482 0.415 0.602 0.481 0.414 0.603 0.333 0.259 0.474 0.360 0.287 0.500
HAHE [23] 0.498 0.426 0.610 0.497 0.429 0.614 0.343 0.270 0.484 0.378 0.306 0.515
UniHR 0.491 0.417 0.618 0.493 0.420 0.621 0.348 0.278 0.482 0.382 0.313 0.513
Link Prediction on HKG.
We compare our method with previous methods on the WD50K and WikiPeople datasets shown in Table 2. We can observe that the models based on transformers (i.e., StarE, GRAN, HyNT, HAHE) demonstrate significantly better performance compared to other models. We attribute this to the transformer’s superior ability to capture long distance dependencies within H-Facts. Among these methods, it can be seen that our proposed UniHR achieves competitive results with the state-of-the-art method HAHE, which means our method effectively captures hierarchical fact information. Compared to the GNN-based method StarE, we achieve improvements of 3.9 points (12.6%) in MRR, 4.4 points (18.8%) in Hits@1 and 3.0 points (6.6%) in Hits@10 on WD50K. This indicates that the performance of StarE’s customized graph neural network is limited by its inability to flexibly capture key-value pair information. Moreover, since the embeddings for newly added fact nodes and relation nodes are computed from atomic facts through simple MLPs, so our training parameters do not significantly increase.
Table 3: Results of base link prediction (left) and triple prediction (right). All baselines’ results are taken from [37]. We pick the best variants of BiVE [7] and NestE [37].
FBH DBHE FBH FBHE DBHE
Model MRR H@10 MRR H@10 MR MRR H@10 MR MRR H@10 MR MRR H@10
Base link prediction Triple prediction
QuatE [45] 0.354 0.581 0.264 0.440 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE [16] 0.356 0.583 0.274 0.446 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP [42] 0.315 0.486 0.233 0.357 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM [12] 0.317 0.490 0.237 0.359 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL [24] 0.310 0.526 0.220 0.364 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE [7] 0.370 0.607 0.274 0.422 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE [37] 0.371 0.608 0.289 0.443 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
Link Prediction on NKG.
Our experiments on the NKGs consist of two tasks: base link prediction and triple prediction. From the results in Table 3, we can see that our proposed UniHR obtains competitive results as the first method to capture global structural information of NKGs. For base link prediction task on triple-based KGs, UniHR achieves considerable improvements. Of particular note, the MRR of FBHE increases by 8.1%.
For triple prediction, we perform best on FBH and DBHE datasets, especially obtaining an improvement of 2.4 points in MRR on FBH, and achieve the second-best performance on FBHE, which suggests that structural information is also valuable for NKG and UniHR can effectively capture the heterogeneity of NKG to enhance node embeddings. In particular, as a unified method, we do not use the customized decoder for triples, while previous state-of-the-art methods do. We will further illustrate the effectiveness of UniHR equipped with other decoders in Appendix B.
Model wikidata12k
MRR H@1 H@3 H@10
ComplEx-N3 [20] 0.248 0.143 - 0.489
HyTE [8] 0.253 0.147 - 0.483
TA-DistMult [13] 0.230 0.130 - 0.461
TeRo [39] 0.299 0.198 0.329 0.507
ATiSE [38] 0.252 0.148 0.288 0.462
TGeomE+ [40] 0.333 0.232 0.361 0.546
HGE [26] 0.290 0.176 0.323 0.514
ECEformer [11] 0.262 0.159 0.255 0.462
UniHR 0.333 0.240 0.367 0.527
Link Prediction on TKG.
As shown in Table 4, we achieve competitive results on the wikidata12k, even surpassing TGeomE+ by 3.4% on Hits@1 and 1.7% on Hits@3. However, existing temporal knowledge graph embedding (KGE) methods, such as TGeomE+, often employ time-aware decoders, which are challenging to generalize to other types of KGs. In contrast, our approach efficiently encodes timestamps as atomic nodes only during initialization and learns temporal information through message passing on graph structure, demonstrating that graph structure information is also beneficial for temporal knowledge graphs, highlighting the effectiveness of our UniHR.
Table 5: Ablation studies on triple prediction task.
Variant FBH FBHE DBHE
MR MRR H@10 MR MRR H@10 MR MRR H@10
w/o initial $\textbf{h}_{f}$ 5.22 0.909 0.980 5.86 0.767 0.885 2.56 0.794 0.978
w/o $\textbf{W}_{r}$ 3.06 0.944 0.992 5.98 0.792 0.885 2.30 0.850 0.976
w/o intra-fact MP 3.97 0.897 0.972 5.26 0.754 0.883 2.02 0.842 0.983
w/o $\omega_{\tau(r)}$ 2.70 0.934 0.992 5.46 0.782 0.888 2.69 0.810 0.973
w/o $\textbf{W}_{\lambda(r)}$ 2.50 0.941 0.992 5.69 0.778 0.889 2.37 0.810 0.978
w/o inter-fact MP 2.61 0.913 0.991 4.56 0.776 0.887 2.11 0.827 0.986
UniHR 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
5.3 Ablation Study on HiSL
We conduct ablation experiments on triple prediction, the most relevant task to fact nodes. As shown in Table 5, all variants with certain modules or parameters removed exhibit a decrease in performance. We can conclude that intra-fact and inter-fact message passing modules both play crucial roles, allowing UniHR to more fully represent the current fact node with enhanced structural information. We also change the initialization of fact node embeddings to learnable embeddings (w/o initial $\textbf{h}_{f}$), and results indicate that initializing fact node representations is essential. It highlights key information in the facts, mitigating the noise introduced by excessive neighbors.
5.4 Potential of Unified Representation
Table 6: Results on hyper-relational TKG datasets.
Model WiKi-hy YAGO-hy
MRR H@1 H@3 H@10 MRR H@1 H@3 H@10
TeRo [39] 0.572 0.473 0.640 0.727 0.760 0.720 0.782 0.822
HGE [26] 0.602 0.507 0.666 0.765 0.790 0.760 0.814 0.837
StarE [12] 0.565 0.491 0.599 0.703 0.765 0.737 0.776 0.820
GRAN [34] 0.661 0.610 0.679 0.750 0.808 0.789 0.817 0.842
HyNT [6] 0.537 0.444 0.587 0.723 0.763 0.724 0.787 0.836
HypeTKG [10] 0.687 0.633 0.710 0.789 0.832 0.817 0.838 0.857
UniHR 0.681 0.618 0.717 0.774 0.823 0.801 0.839 0.859
Generalize to Hyper-relational TKGs.
Due to the unity of representation, UniHR can easily generalize to compositional types of KGs, like hyper-relational TKGs (HTKGs) [10] that integrate HKGs and TKGs. Every H-fact in the HTKG is coupled with a timestamp explicitly indicating its time validity. From results in Table 6, UniHR offers a significant performance improvement in link prediction tasks on hyper-relational TKGs compared to TKG-specific and HKG-specific models, demonstrating a good collaborative modeling ability for auxiliary key-value pairs and timestamps. Our UniHR even achieves competitive performance with the specialized hyper-relational TKG model HypeTKG without complex module stacking.
Table 7: Results of separate and joint training on NKG.
FBH DBHE FBH DBHE
Model MRR H@10 MRR H@10 MRR H@10 MRR H@10
Base link prediction Triple prediction
UniHR 0.401 0.619 0.296 0.448 0.946 0.993 0.862 0.987
UniHRJoint 0.402 0.632 0.298 0.452 0.948 0.993 0.857 0.989
Joint Learning on Different Tasks of KGs.
For link prediction on NKGs, the two subtasks, namely base link prediction and triple prediction, share the same graph during the message-passing phase under our unified representation form. Therefore, we attempt joint training on the two tasks using the NKG dataset, as shown in Table 7. We observe that the results of joint training are generally superior to those of separate training, indicating that nested facts and atomic facts can mutually enhance and supplement the semantics.
Table 8: Results of separate training and joint training on the HKG and TKG dataset, where identical entities and relations share the same embeddings. WikiPeople- and wikidata12k- represent the filtered test sets.
WikiPeople- wikidata12k-
Model subject/object all entities subject/object
MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10
UniHRJoint 692.7 0.488 0.409 0.533 0.629 686.5 0.490 0.414 0.536 0.632 489.5 0.315 0.222 0.346 0.498
Joint Learning on Different Types of KGs.
We suppose that unified representation makes it possible to develop pre-trained models that integrate multiple types of KGs. To explore its potential, we conduct joint learning on different types of KGs. Therefore, we construct a hybrid dataset called wikimix which includes two subsets of Wikidata [33], namely HKG dataset WikiPeople and TKG dataset wikidata12k, which encompass 3547 identical entities and 18 identical relations. Due to the different types of facts, there are no identical facts in these two subsets. To further prevent data leakage, we filter out 537 entries from the wikidata12k test set whose main triples appear in the H-Facts of WikiPeople train set, and 384 entries from the WikiPeople test set whose main triples appear in wikidata12k train set. Statics of wikimix are given in Table 1.
From the results in Table 8, it is evident that joint learning outperforms separate learning across most metrics. Notably, there are improvements of 17.1% and 40.2% in MR metric on wikipeople- and wikidata12k- datasets, respectively. This indicates that more complex structural interactions and diverse types of training data are beneficial. Moreover, our UniHR demonstrates good scalability and effectiveness in integrating multiple types of KGs.
Table 9: Number of training parameters and times.
Dataset Model Training Params Training Times
WD50K StarE [12] 10.81M 48h+
HyNT [6] 29.61M 22.3h
HAHE [23] 42.35M 29.2h
UniHR 10.55M 23.4h
wikidata12k ATiSE [38] 31.46M 5.8h
HGE [26] 3.44M -
ECEformer [11] 11.99M 5.6h
UniHR 3.68M 3.6h
YAGO-hy HypeTKG [10] 10.83M 37.5h
UniHR 4.37M 17.6h
5.5 Efficiency Analysis
Compared to existing state-of-the-art methods, UniHR does not significantly increase the overhead in terms of model training parameters and runtime as shown in Figure 9. The embeddings of relation nodes and fact nodes are computed using a simple MLP based on atomic nodes and relations, which does not significantly increase the number of training parameters. For runtime, we do not use the whole graph during all message passing phases and adopt a dropout strategy to prevent overfitting, which can effectively enhance training efficiency. Therefore, our UniHR achieves a superior trade-off between model effectiveness and efficiency.
6 Conclusion and Future Work
In this paper, we propose UniHR, a unified hierarchical knowledge graph representation learning framework consisting of a Hierarchical Data Representation (HiDR) module and a Hierarchical Structure Learning (HiSL) module. The HiDR form unifies the hyper-relational facts, nested facts and temporal facts into the form of triples, overcoming the limitations of customized encoders for different forms of facts. Moreover, HiSL captures local semantic information within facts and global structural information between facts. Our UniHR achieves the best or competitive performance across five types of KGs, and highlight the strong potential of unified representations.
In the future, we believe that the unification of knowledge graph representation learning methods is a developing trend that makes it possible to develop unified pre-trained models based on multiple types of KGs. Moreover, we will investigate how to generalize our HiDR form to more complex tasks such as inductive reasoning [30] and multi-modal scenarios, etc.
Supplemental Material Statement:
The dataset, code, and hyperparameter details are available at https://github.com/Lza12a/UniHR.
Appendix
Appendix 0.A Related Works
0.A.1 Link Prediction on Triple-based KGs
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models [4, 29] have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE [4], RotatE [29], and ConvE [9].
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) [31, 25, 41] proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN [31] aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN [41] learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are hard to generalize to beyond-triple KGs.
Table 10: Base link prediction on FBHE, FBH and DBHE. All baselines’ results are taken from [37]. The best results among all models are written bold, while the second are underlined. The $s_{f}$ and $s_{h}$ denote $\left(f,has\;head\;entity,h\right)\ \left(f,has\;tail\;entity,t\right)$ and $\left(h,r,t\right)$ two types of scoring method respectively. For BiVE [7] and NestE [37], we pick their variants with best performance.
Model FBHE/FBH DBHE
MRR Hits@10 MRR Hits@10
QuatE 0.354 0.581 0.264 0.440
BiQUE 0.356 0.583 0.274 0.446
Neural-LP 0.315 0.486 0.233 0.357
DRUM 0.317 0.490 0.237 0.359
AnyBURL 0.310 0.526 0.220 0.364
BiVE 0.370 0.607 0.274 0.422
NestE 0.371 0.608 0.289 0.443
UniHR + ConvE $s_{h}$ 0.397 0.622 0.289 0.443
UniHR + ConvE $s_{f}$ 0.375 0.596 0.307 0.471
UniHR + Transformer 0.401 0.619 0.296 0.448
Appendix 0.B Decoder Analysis
To explore the effectiveness of our UniHR encoding further, we pair UniHR with different decoders and evaluated them on triple prediction task. In addition to the previously mentioned unified framework UniHR + Transformer, we also experiment on UniHR + ConvE with two scoring strategies. The ConvE [9] is the decoder customized for triples and its scoring function is $vec\left(\sigma\left(\left[\mathbf{\tilde{h}}_{h};\mathbf{\tilde{e}}_{r}\right]*\psi\right)\right)$, where $\mathbf{\tilde{h}}_{h}$ and $\mathbf{\tilde{e}}_{r}$ represent reshaped 2D embeddings of head entity $h$ and relation $r$, and * is a convolution operator. The $vec\left(\cdot\right)$ and $\psi$ are denoted as the vectorization function and a set of convolution kernels.
Due to our special representation, there exists two scoring methods for atomic triples, thus we present the base link prediction results separately for each scoring method. The $s_{f}$ represents scoring triples $\left(f,has\;head\;entity,h\right)$and $\left(f,has\;tail\;entity,t\right)$, and $s_{t}$ represents scoring $\left(h,r,t\right)$. The performance of base link prediction is shown in Table 10. Notably, FBH and FBHE share identical atomic facts, resulting in the same performance. It can be observed that regardless of the scoring method employed, we both achieve competitive performance, especially with scoring $\left(h,r,t\right)$ on FBH and scoring $\left(f,has\;head\;entity,h\right)\,\left(f,has\;tail\;entity,t\right)$ on DBHE. We attribute the differences in performance under different scoring methods to dataset characteristics. The DBHE dataset is relatively smaller, and scoring method $s_{f}$ effectively alleviates overfitting problem. Conversely, for larger datasets FBH, scoring based on $\left(h,r,t\right)$ minimizes information loss.
Table 11: Triple prediction on FBHE, FBH and DBHE. All baselines’ results are taken from [37]. The best results among all models are written bold. For BiVE [7] and NestE [37], we pick their variants with best performance.
Model FBH FBHE DBHE
MR MRR Hits@10 MR MRR Hits@10 MR MRR Hits@10
QuatE [45] 145603.8 0.103 0.114 94684.4 0.101 0.209 26485.0 0.157 0.179
BiQUE [16] 81687.5 0.104 0.115 61015.2 0.135 0.205 19079.4 0.163 0.185
Neural-LP [42] 115016.6 0.070 0.073 90000.4 0.238 0.274 21130.5 0.170 0.209
DRUM [28] 115016.6 0.069 0.073 90000.3 0.261 0.274 21130.5 0.166 0.209
AnyBURL [24] 108079.6 0.096 0.108 83136.8 0.191 0.252 20530.8 0.177 0.214
BiVE [7] 6.20 0.855 0.941 8.35 0.711 0.866 3.63 0.687 0.958
NestE [37] 3.34 0.922 0.982 3.05 0.851 0.962 2.07 0.862 0.984
UniHR + Transformer 2.46 0.946 0.993 5.20 0.793 0.890 1.90 0.862 0.987
UniHR + ConvE 3.00 0.900 0.983 6.27 0.909 0.986 2.06 0.876 0.978
Table 11 shows the results of triple prediction on three benchmark datasets. Among all baselines, Quate, Bique, Neural-LP, Drum, and AnyBURL struggle to model the mapping relationship between atomic facts and nested facts. Furthermore, prior works [7] do not guarantee that all atomic facts in the nested fact test set are present in the training set as entities, which shifts the problem from a transductive setting to an inductive setting, leading to significant performance gaps between these baselines. On most metrics, our method outperforms BiVE and NestE which are specifically modeled for nested facts. Notably, NestE fully preserves the semantics of atomic facts. However, on the FBHE dataset, UniHR + ConvE achieves an improvement of 0.58 (6.4%) points in MRR and 0.24 (2.4%) points in Hits@10 compared to the state-of-the-art model NestE and the second-best performance after UniHR + Transformer on the FBH and DBHE datasets, demonstrating UniHR’s powerful graph structure encoding capabilities. We also carry out ablation experiments on UniHR + ConvE as shown in Table 11. Performance declines are observed after removing any part of the HiSL module, showing the significance of HiSL for hierarchical encoding.
@@ Appendix 0.B Decoder Analysis
Appendix 0.C Hyperparameter Settings

### v4

@@ Appendix C B More Ablation Study
Table 9: The statistics of diverse knowledge graphs dataset, where “with Q(%)” and “Arity” column respectively denote the number of facts with auxiliary key-value pairs and the range of arity of hyper-relational facts, “N-Fact” is the number of nested fact, “N-Rela” is the number of nested relation, the “AF(%)” column denotes the number of atomic facts in nested facts.
@@
WikiPeople wikidata12k
Model subject/object all entities subject/object
MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10 MR MRR H@1 H@3 H@10
UniHRJoint 692.7 0.489 0.409 0.530 0.629 686.5 0.493 0.418 0.536 0.632 493.6 0.317 0.224 0.348 0.503
Table 11: Results of separate training and joint training on hybrid KG dataset wikimix, where identical entities and relations share the same embeddings. WikiPeople and wikidata12k represent the filtered test sets.
FBH DBHE FBH DBHE
Model MRR H@10 MRR H@10 MRR H@10 MRR H@10
Base link prediction Triple prediction
UniHR 0.401 0.619 0.296 0.448 0.946 0.993 0.862 0.987
UniHRJoint 0.404 0.632 0.298 0.454 0.949 0.994 0.860 0.989
Table 12: Results of separate and joint training on NKG.
Appendix D C More Related Works
Link Prediction on Triple-based KGs.
Most existing techniques in KG representation learning are proposed for triple-based KGs. Among these techniques, knowledge graph embedding (KGE) models (Bordes et al. 2013; Sun et al. 2018) have received extensive attention due to their effectiveness and simplicity. The idea is to project entities and relations in the KG to low-dimensional vector spaces, utilizing KGE scoring functions to measure the plausibility of triples in the embedding space. Typical methods include TransE (Bordes et al. 2013), RotatE (Sun et al. 2018), and ConvE (Dettmers et al. 2018).
Depending on the KGE model alone has limitation of capturing complex graph structures, whereas augmenting global structural information with a graph neural network (GNN) (Vashishth et al. 2019; Nathani et al. 2019; Xu, Bao, and Liu 2023) proves to be an effective approach for enhancement. The paradigm of combining GNN as encoder with KGE scoring function as decoder helps to enhance the performance of KGE scoring function. These GNN methods design elaborate message passing mechanisms to capture the global structural features. Typically, CompGCN (Vashishth et al. 2019) aggregates the joint embedding of entities and relations in the neighborhood via a parameter-efficient way and MA-GNN (Xu, Bao, and Liu 2023) learns global-local structural information based on multi-attention. These methods achieve impressive results on triple-based KGs but are challenging to generalize to beyond-triple KGs. Additionally, these GNNs usually focus more on modeling global information while neglecting local information.
Why We Need Unified Representation?
Firstly, real-world data sources naturally contain heterogeneous data forms (Liu et al. 2025b). For example, Wikidata (Vrandečić and Krötzsch 2014) includes triple-based facts, hyper-relational facts, temporal facts, and even nested facts, financial reports and research papers usually contain various modalities such as tables, images, and text. Unified representation aims to integrate multiple heterogeneous data within a single semantic structure, avoiding the need for cross-type data conversion or independent processing during use. In many real-world scenarios (e.g., question answering, recommendation, or transfer learning), the research community is constantly exploring the benefits of unified representations. For question answering (QA), TrustUQA (Zhang et al. 2025b) demonstrates a unified representation allows the knowledge base to contain a broader range of knowledge types, thereby simplifying the burden of the knowledge retrieval module and enabling the QA system to cover a wider variety of questions and answers. For recommendation systems, DURation (Lu et al. 2022) with unified representation effectively eliminate information redundancy caused by separate module designs and help capture cross-dimensional relational knowledge, thus improving the comprehensiveness and accuracy of recommendations. Moreover, UniHR not only meets the requirements of real-world KG reasoning tasks (involving multiple types of facts), but we also believe that unified representation itself is beneficial for transfer learning. It not only facilitates transfer learning across different types of KGs, but also enables pre-training across multiple KG types (as demonstrated in Section 5.4). This further promotes more effective pre-training and transfer learning in the field of knowledge graphs.
Appendix E D Detailed Results of Our “Potential”
Table 11 shows the results of “Joint Learning on Different Tasks of KGs” section. Table 12 shows the results of “Joint Learning on Different Types of KGs” section.
Appendix F E Dataset Details
Table 9 shows the details of the two hyper-relational knowledge graph (HKG) benchmark datasets: WikiPeople, WD50K three nested factual knowledge graph (NKG) benchmark datasets: FBH, FBHE, DBHE, and the temporal knowledge graph (TKG) benchmark dataset wikidata12k.
Among them, WikiPeople is a dataset derived from Wikidata (Vrandečić and Krötzsch 2014) concerning entities type “human”. WikiPeople filter out the elements which have at least 30 mentions as key-value pairs. WD50K is a high-quality dataset extracting from Wikidata statements and avoiding the potential data leakage which allows triple-based models to memorize main fact in the H-Facts of test set. The “with Q(%)” column in Table 9 denote the number of facts with auxiliary key-value pairs and the “Arity” column denote range of the number of entities in hyper-relational facts.
Nested factual knowledge graph datasets FBH and FBHE (Chung and Whang 2023) are constructed based on FB15k237 from Freebase (Bollacker et al. 2008) while DBHE is based on DB15K from DBpedia (Lehmann et al. 2015). FBH contains nested facts that can be only inferred inside the atomic facts, while FBHE and DBHE contain externally-sourced nested relation crawling from Wikipedia articles, e.g., NextAlmaMater and SucceededBy.
Temporal knowledge graph dataset wikidata12K is also a subset of Wikidata (Vrandečić and Krötzsch 2014), which represents the time information $\tau\in\mathcal{T}$ as time intervals.
Appendix G F Hyperparameter Settings